OpenStudy (anonymous):

The sum of two consecutive even integers is –46. What are the numbers? (Points : 1) –26 and –20 –22 and –20 –26 and –24 –24 and –22

OpenStudy (anonymous):

@IMStuck @iGreen @LieutenantGeneral

OpenStudy (igreen):

-26 and -20 = -46 -22 + -20 = -42 -26 + -24 = -50 -24 + -22 = -46

OpenStudy (igreen):

Can you choose more than one option..?

OpenStudy (lieutenantgeneral):

-24 and -22

OpenStudy (anonymous):

So A or D, and no

OpenStudy (anonymous):

I was thinking D too

OpenStudy (lieutenantgeneral):

yep you add them

OpenStudy (anonymous):

okay :) thanks both of you :)

OpenStudy (lieutenantgeneral):

you welcome

OpenStudy (anonymous):

i dont know who to give the medal too

OpenStudy (anonymous):

I want to give it to both ^_^

OpenStudy (anonymous):

YESSS its D :)))))

OpenStudy (igreen):

I already gave one to @LieutenantGeneral

OpenStudy (anonymous):

thanks ^